if the adjacent side of a parallelogram are 36cm and 20cm.  the distance between the longer side is 12cm find the distance between the shorter side.

Grade:7

1 Answers

Naman Agarwal
7 years ago
Area of parallelogram is Base*heightSo,we can take the area of parallelogram from two ways.Base1 * Height1 =Base2 * Height236 * 12 = 20 * Height2Height2= 36*12/20Height2 = 21.6cm